OpenStudy (anonymous):

ABCD is a parallelogram. If measurement D= 85, then what is measurement C?

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

D=B=85 D+B=170 360-170=190 A=C=190/2 =95 degree